OpenStudy (katherinesmith):

What is the equation of a parabola with vertex (0, 0) and a focus at (2, 0)

OpenStudy (anonymous):

didn't we just do this?

OpenStudy (katherinesmith):

nope, we did (0, 0) and (0, 2)

OpenStudy (katherinesmith):

(0, 6)*

OpenStudy (anonymous):

oh i see, this time the focus is on the \(x\) axis looks like 

OpenStudy (katherinesmith):

i could've used the equation you gave me but didn't know if something different happened because the x and y's are switched

OpenStudy (anonymous):

so this time instead of \(4py=x^2\) we have \(4px=y^2\)

OpenStudy (anonymous):

yeah if it opens left or right, the \(y\) term is squared, not the \(x\) term

OpenStudy (anonymous):

but now it is easy right? you have \[4px=y^2\] and \[p=2\]

OpenStudy (katherinesmith):

8x = y^2

OpenStudy (katherinesmith):

what about something like the equation of a parabola with vertex (0, 0) and directrix x = -3